OpenStudy (anonymous):

Solve this system of linear equations. 4x+5y=-24 16x-2y=-8

OpenStudy (bahrom7893):

do what we have been doing, eliminate and put it into the other equation..

OpenStudy (anonymous):

Okay, but idk how you made the equation, it confuses me :(

OpenStudy (bahrom7893):

Okay here: 4x+5y=-24 Slove for y in terms of x

OpenStudy (amistre64):

equations can be manipulated to bend to our will; we can multiply them by any number and they retain their inherent value/ proportions: Lets choose a number that will benefit us; lets see what happens when we multiply the top equation here by -4 4x+5y=-24 ; *-4 16x-2y=-8 -16x-20y=96 16x- 2y=-8 ; now add the results together -------------- 0x -22y =88 ; now we solve for y

OpenStudy (anonymous):

x=-1 y =-4